jsinrod:

Grace goes out to lunch. The bill, before tax and tip, was $11.10. A sales tax of 7% was added on. Grace tipped 16% on the amount after the sales tax was added. How much was the sales tax? Round to the nearest cent.

jsinrod:

@hero

jsinrod:

this is confusing

Hero:

7% just means \(\dfrac{7}{100}\) So to calculate the amount of the bill after ADDING sales tax, it's BILL\(+\) BILL \(\times\) SALES TAX

Hero:

Actually, they only want the sales tax so do this instead.

Hero:

SALES TAX = PERCENTAGE \(\times\) BILL

Hero:

So: SALES TAX = \(\dfrac{7}{100} \times \left(11 + \dfrac{10}{100}\right)\)

Hero:

You can also simply use a calculator to find the answer: Sales Tax = (0.07)(11.10)